OpenStudy (anonymous):

solve the equation -7(2x-1)=35 x=???

OpenStudy (hba):

-14x+7=35 -14x=28 x=-2 answer

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

THANK YOU GUYS!!!!!!!

OpenStudy (hba):

give me best response @Gerber96

OpenStudy (anonymous):

Yasir is right

OpenStudy (anonymous):

hba is right too lol

OpenStudy (anonymous):

thanx timo. i am right but gerber did well..:-)